# AgMMU: A Comprehensive Agricultural Multimodal Understanding and Reasoning Benchmark

### v1

@@ A Evaluated Large Vision Language Models
Our evaluation and analysis are conducted mainly on the group of models listed in Table 2 in the main paper. We have chosen models such that they cover most of the popular and best-performing methods used by recent multimodal understanding and reasoning work. In this part, we discuss all the models we have used in our experiments and explain their evaluation details, the public checkpoints we have chosen, and display the prompts we used to adapt the model to our datasets.
During evaluation, we chose to follow the standard prompt provided by the authors whenever possible for multiple-choice and short-answer questions. When the prompt is not provided for the model, we select a custom prompt that is created through several iterations of prompt engineering to select the one that produces the most effective results. For all open-source models, we only chose to only include the first associated image to each question even though some models support multi-image inference. The image is always included as the prefix. This way, we remain consistent and can reproduce the results.
Proprietary Models  . We used three proprietary models in our evaluation: GPT-4o [1], Gemini 1.5 Pro [33], and Claude 3 Haiku [40]. Below we note the model version and access date used for evaluation.
  • GPT-4o: November 1-13, 2024.
@@
BioCLIP   [38]. BioCLIP is trained on Tree-Of-Life 10M, a large biological species dataset that allows it to learn hierarchical species names and relationships. By learning the hierarchical species relationship, it has greater generalization to species that were not in the training data, but closely related. We use the official implementation and checkpoint [^1] https://huggingface.co/imageomics/bioclip with a CLIP ViT-B/16 base model in our evaluation. During evaluation, we follow the evaluation methods of the original paper. Additionally, we only evaluated this model on Disease/Issue Identification, Insect/Pest Identification, and Species Recognition categories due to its major focus of species classification.
@@
LLaVA-OneVision   [21]. LLaVA-OneVision builds on LLaVA-NeXT with the capability to analyze single images, multi-images, and video scenarios. Most impressively, it allows for video understanding through task transfer from images but this is not really explored in our evaluations. We use the official implementation and checkpoint [^6] https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf that uses a base architecture consisting of SigLIP-SO400M-Patch14-384 and Qwen2-7b in our evaluation.
LLaVA-1.5-7B / LLaVA-1.5-13B   [24]. LLaVA introduces the idea of instruction tuning a multimodal model with GPT-4 generated instruction-following data for associated images. This gives it the ability to achieve impressive abilities to act as an instruction-following general agent. We use the official implementation and checkpoint [^7] https://huggingface.co/llava-hf/llava-1.5-7b-hf [^8] https://huggingface.co/llava-hf/llava-1.5-13b-hf with a CLIP ViT-L/14 vision backbone and Vicuna1.5-7B / Vicuna1.5-13B in our evaluation.
@@ B Dataset Curation Details
This section outlines the three-stage curation pipeline of AgMMU and describes the prompts designed for each question type and subdomain.
Figure A: Prompts for our three-stage data curation pipeline.
Figure B: Examples included in prompt during agriculture domain categorization (step 1)
Figure C: Examples included in prompt during knowledge extraction (step 2) based on agriculture subdomain type
@@
Figure E: Prompts for our LLM as judge on short-answer OEQ.
Figure F: Prompts for the answer splitting step of our LLM as judge on multi-sentence (long-answer) OEQ.
Figure G: Prompts for the evaluation step of our LLM as judge on multi-sentence (long-answer) OEQ.
@@ B.3 Stage 3: Question Generation
This contextualization ensures that the generated questions integrate both user-provided information and extracted context seamlessly. The result is a set of comprehensive and ”fair” evaluative questions, designed to effectively assess multimodal agricultural understanding and reasoning.
Figure H: GPT-4o accuracy with increasing MCQ options. Model performance on MCQs across different categories, comparing accuracy scores when varying the number of answer options (4, 5, and 6). We observe a 5-10% difference in accuracy across categories between the 4-option and 6-option configurations, with performance generally decreasing as the number of options increases.
C More Evaluation and Design Choices
LLM as Judge.   To perform evaluation on few-word and multi-statement OEQ responses, we implemented the LLM-as-judge methodology using Qwen-7B. Our prompts for few-word responses (Figure E) and multi-statement responses (Figure F, G) contain several in-context examples to guide the LLM to correctly categorize the answer as “correct”, “incorrect”, “partially correct”, and “irrelevant”.
@@ C More Evaluation and Design Choices
The results, shown in Figure  H, indicate that accuracy decreases as the number of answer choices increases. Specifically, we observed a 5-10% reduction in accuracy between the 4-option and 6-option configurations. This trend suggests that the model might rely on a process of elimination when selecting an answer, making it more challenging to identify the correct response as the number of options increases. While the decrease in accuracy is not overly significant, we think it justifies our choice to use four options for MCQs.
@@ D More Dataset Visualization
In Table I, we demonstrate more samples in AgMMU with questions and multiple choice answers.
Figure I: Additional visualization of samples in AgMMU. Ground truth selections of each question are highlighted in yellow.
@@ E Limitations and Future Work
While our work makes significant contributions to agricultural benchmark development and MLLM evaluation through knowledge-intensive tasks, we acknowledge several limitations and identify promising directions for future research in this section.
Advanced Utilization of Training Data.   Although our curated dataset, AgBase-200K, has proven significant effectiveness for fine-tuning MLLMs [24] as shown in Section 4 and Figure 7, its potential extends beyond our current usage. As a comprehensive knowledge repository, the dataset presents opportunities for knowledge retrieval and augmented generation approaches. In particular, the development of vision-centric multimodal retrieval augmented generation (RAG) systems remains an under-explored yet promising direction. This alternative could enable more effective knowledge extraction and utilization from our dataset, potentially improving model performance on agricultural understanding tasks. We leave the exploration of these advanced techniques for future work.
Expanded Model Coverage and Evaluation Protocols.   While our current study encompasses several state-of-the-art and most commonly used MLLMs for zero-shot evaluation and fine-tuning analysis, we acknowledge that these represent only a subset of available multimodal architectures and methodologies. To enhance the robustness and generalizability of our findings, we plan to incorporate a broader spectrum of MLLMs. Additionally, we plan to conduct more extensive ablation studies and comparative analyses across different model scales and architectures. This comprehensive evaluation will provide deeper insights into the relative strengths and limitations of various approaches in agricultural understanding tasks.
@@ F Societal Impact
We anticipate no negative societal impact of our work. Our dataset is ethically designed, respecting the privacy of Extension.org users by removing personal identifying information such as name, gender, and location. Additionally, we have verified to the best of our ability to ensure the removal of images that contain human faces. During dataset curation, we put in great effort to eliminate bias by creating a dataset representative of the original Extension.org questions as well as a balanced dataset across all question types.

### v2

@@ A Evaluated Large Vision Language Models
Our evaluation and analysis are conducted mainly on the group of models listed in Table 2 in the main paper. We have chosen models such that they cover most of the popular and best-performing methods used by recent multimodal understanding work. In this part, we discuss all the models we have used in our experiments and explain their evaluation details, the public checkpoints we have chosen, and display the prompts we used to adapt the model to our datasets.
During evaluation, we chose to follow the standard prompt provided by the authors whenever possible for multiple-choice and short-answer questions. When the prompt is not provided for the model, we select a custom prompt that is created through several iterations of prompt engineering to select the one that produces the most effective results. The images are always included as the prefix.
Proprietary Models.   We used three proprietary models in our evaluation: GPT-o4-mini [1], Gemini 1.5 Pro [34], and Claude 3 Haiku [41]. Below we note the model API version used for evaluation.
  • GPT-o4-mini: May 13-15, 2025.
@@
LLaVA-OneVision   [22]. LLaVA-OneVision builds on LLaVA-NeXT with the capability to analyze single images, multi-images, and video scenarios. Most impressively, it allows for video understanding through task transfer from images but this is not explored in our evaluation. We use the official implementation and checkpoint [^5] https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf that uses a base architecture consisting of SigLIP-SO400M-Patch14-384 and Qwen2-7b in our evaluation.
LLaVA-1.5-7B / LLaVA-1.5-13B   [25]. LLaVA introduces the idea of instruction tuning a multimodal model with GPT-4 generated instruction-following data for associated images. This gives it the ability to achieve impressive abilities to act as an instruction-following general agent. We use the official implementation and checkpoints [^6] https://huggingface.co/llava-hf/llava-1.5-7b-hf [^7] https://huggingface.co/llava-hf/llava-1.5-13b-hf with a CLIP ViT-L/14 vision backbone and Vicuna1.5-7B / Vicuna1.5-13B in our evaluation.
@@ B Dataset Curation Details
This section outlines the multi-stage curation pipeline of AgMMU and describes the prompts designed for each question type and subdomain.
Figure A: Prompts used in different stages of our data curation pipeline.
Figure B: Examples included in prompt during the agriculture domain categorization (step 1).
Figure C: Examples included in prompt during the knowledge extraction (step 2) based on agriculture subdomain type.
@@
Figure E: Grading prompt for our LLM-as-judge on short-answer OEQ.
Figure F: Unique examples included for short-answer categories added to the grading prompt for our LLM-as-judge on short answer OEQ.
Figure G: Prompt for categorizing statements in our LLM-as-judge on multi-sentence (long-answer) OEQ.
Figure H: Unique examples included for multi-statement categories added to the grading prompt for our LLM-as-judge on multi-sentence (long-answer) OEQ.
@@ B.3 Stage 3: Question Generation
This contextualization ensures that the generated questions integrate both user-provided information and extracted context seamlessly. The result is a set of comprehensive and “fair” evaluative questions, designed to effectively assess multimodal agricultural understanding.
Figure I: GPT-4o accuracy with increasing MCQ options. Model performance on MCQs across different categories, comparing accuracy scores when varying the number of answer options (4, 5, and 6). We observe a 5-10% difference in accuracy across categories between the 4-option and 6-option configurations, with performance generally decreasing as the number of options increases.
B.4 Final Stage: Human Verification
To guarantee the quality of the evaluation questions, we implemented a human verification process that validates faithfulness, certainty, quality, and MCQ feasibility. The data was distributed through an HTML file containing AgMMU questions and answers, original user questions, expert answers, and corresponding images. Each annotator was given a corresponding Excel file where the user just has to mark false (uncheck the box) for each condition not met per question. To further assist the annotator, we provided a few complex examples of questions that meet and do not meet the requirements, functioning as in-context examples. After collecting these data, only the completely unproblematic ones (all boxes remain checked) were kept.
Faithfulness: Do you think the question, ground-truth, and context extract faithful information from the original farmer question? Our questions are directly based on the original questions and this step functions as a sanity check ensuring the quality of our dataset. The annotator needs to read through the question and the original conversations between the user and the expert.
Certainty: Is the expert certain about the answer? As our ground truth answers are extracted from the expert answers, we only want to include those that are very certain. A higher certainty from the expert means that it is more likely to be correct. We observe that the behaviors of the annotator are to read the responses from the expert and look for keywords like “may,” “not sure,” “you have to go to a lab for further inspection.”
Quality: Are the images suitable for answering the questions? (Images are not in low-resolution, blurs, pure blackness, etc.) Are the image/symptom descriptions visible in the presented images? As our benchmark attempts to evaluate the visual understanding of the models, our human verification removes the questions that do not depend on the images and those with broken images. For example, it is not fair for our question to ask about the fruit of the plant when the submitted photos only capture the leaves of the tree or the image is blurry.
Feasibility: Are all of the wrong choices wrong? The incorrect choices were generated with GPT-4o, so we need to check to ensure there are no multiple correct answers or an answer that overlaps with the correct answer and remove. For example, the wrong choice might be the common name of a species displayed by the scientific name in the ground truth.
C More Evaluation, Implementation, and Design Choices
Figure J: Evaluation scores across five error categories for three fine-tuning setups using the AgBase dataset. Models are fine-tuned on: (1) LLaVA 10k SFT—a mix of AgBase and 10k LLaVA-Instruct samples, (2) LLaVA 57k SFT—a 50-50 blend of AgBase and LLaVA’s original 57k SFT samples, and (3) LLaVA Species SFT—a specialized set focused on species identification with contextual augmentation.
LLM-as-judge.   To perform evaluation on few-word and multi-statement OEQ responses, we implemented the LLM-as-judge methodology using GPT-4.1. Our prompts for few-word responses (Figure EF) and multi-statement responses (Figure G, H) contain several in-context examples based on the question category to guide the LLM to correctly categorize the answer as “correct,” “incorrect,” “partially correct,” and “irrelevant.”
@@ C More Evaluation, Implementation, and Design Choices
The results, shown in Figure  I, indicate that accuracy decreases as the number of answer choices increases. Specifically, we observed a 5-10% reduction in accuracy between the 4-option and 6-option configurations. This trend suggests that the model might rely on a process of elimination when selecting an answer, making it more challenging to identify the correct response as the number of options increases. While the decrease in accuracy is not overly significant, we think it justifies our choice to use four options for MCQs.
Implementation of AgBase fine-tuning.   We fine-tune the LLaVA-v1.5-7B model using a LoRA-based setup. The training is performed with a learning rate of 2e-4, without weight decay, and a cosine learning rate schedule with a 3% warm-up ratio. We use a per-device batch size of 16 with gradient accumulation steps set to 2, resulting in an effective batch size of 64. The model is trained over 2 epochs using 2 NVIDIA A6000 GPUs.
Dataset preparation involved curating structured multi-turn conversations from a horticultural FAQ knowledge base, paired with user-uploaded images. From an initial pool of 367,331 QA-image pairs, we filtered out questions that had a species value in [tree, bee, shrub, weed, wasp, plant, insect, grass, none, moth, beetle, snake, caterpillar, spider, ant, mushroom, fungus], because we observe that questions with these common non-species species extractions often contain vague or uncertain examples. This gives us a high-quality dataset of 57,079 samples. Considering the influence of data mixture for training VLMs, we conduct three fine-tuning experiments. (1) The first experiment involves fine-tuning on a combination of our domain-specific dataset, AgBase, and 10,000 samples from LLaVA’s original instruction-tuning dataset, LLaVA-Instruct-150K. (2) The second experiment employs a 50-50 mixture of AgBase by using 57,079 samples from LLaVA’s original SFT set [25] (3) The third experiment focuses solely on species identification and consists of 18,109 QA pairs constructed by prepending the full original user queries to 33,777 generic identification samples, allowing us to test the effect of user context on classification accuracy.
In Figure J, we find that the LLaVA 10k SFT model achieves a slightly higher overall accuracy (0.25) compared to the LLaVA 57k SFT model (0.24), suggesting that a smaller, well-curated dataset mixed with domain-specific data may be more effective than a larger, more generic one for knowledge-intensive domain fine-tuning. Additionally, the LLaVA Species SFT model, which includes added user query context for species identification, performs worse than the other models in the species category , indicating that this additional context provides limited benefit for classification accuracy.
@@ D More Dataset Visualization
In Figure K, we demonstrate more samples in AgMMU with questions and multiple choice answers.
In Figure L, we demonstrate more samples in AgMMU with open-ended questions and responses. We especially emphasize the long-form responses required from the model for symptom description and management instructions, normally containing multiple facts.
Figure K: Additional visualization of samples in AgMMU. Ground truth selections of each question are highlighted in yellow.
Figure L: Additional visualization of OEQ samples in AgMMU.
@@ E Limitations and Future Work
While our work makes unique contributions to agricultural benchmark development and VLM evaluation through knowledge-intensive tasks, we acknowledge several limitations and identify promising directions for future research in this section.
Advanced Utilization of Training Data.   Although our curated dataset, AgBase, has proven significant effectiveness for fine-tuning VLMs [25] as shown in Section 4 and Figure 6, its potential extends beyond our current usage. As a comprehensive knowledge repository, the dataset presents opportunities for knowledge retrieval and augmented generation (RAG) approaches [5]. In particular, the development of vision-centric multimodal RAG systems remains an under-explored yet promising direction. This alternative could enable more effective knowledge extraction and utilization from our dataset, potentially improving model performance on agricultural understanding tasks. We leave the exploration of these advanced techniques for future work.
Expanded Model Coverage and Evaluation Protocols.   While our current study encompasses several state-of-the-art and most commonly used VLMs for zero-shot evaluation and fine-tuning analysis, we acknowledge that they represent only a subset of available multimodal architectures and methodologies. To enhance the robustness and generalizability of our findings, we plan to incorporate a broader spectrum of VLMs. Additionally, we plan to conduct more extensive ablation studies and comparative analyses across different model scales and architectures. This comprehensive evaluation will provide deeper insights into the relative strengths and limitations of various approaches in agricultural understanding tasks.
@@ F Societal Impact
We anticipate no direct negative societal impact of our work. Our dataset is ethically designed, respecting the privacy of Extension.org users by removing personal identifying information such as name, gender, username, and location. Additionally, we have verified to the best of our ability to ensure the removal of images that contain human faces. During dataset curation, we put in great effort to eliminate bias by creating a dataset representative of the original Extension.org questions as well as a balanced dataset across all question types.